iyoni124:

A mechanic repairs Mr. Estes’ car. The amount for the parts is $48.00 and the rate for the mechanism is $40.00 per hour. Write and solve an equation to find the total amount of repairs to Mr. Estes’ car if the mechanic works for 1.5 hours.

supie:

This is simple, all you have to do it \[y = 48 + 40(1.5)\]\(y=48+40(\frac{15}{10})\) \(y=48+40(\frac{3(5)}{2(5)})\)\[y=48+60\]\(\sf\ and\ 48+60\ equals\ 108 (48+60=108)\ which\ means\ y=108\)